OpenStudy (anonymous):

√5(√5-√3) simplest form ?

OpenStudy (solomonzelman):

\(\large\color{slate}{ \sqrt{5} (\sqrt{5} -\sqrt{3} ) }\). You know that \(\large\color{slate}{ a(b-c)=ab-ac }\)

OpenStudy (solomonzelman):

use that same thing here, but remember that \(\large\color{slate}{ \sqrt{a} \cdot \sqrt{a} =a }\) (for any number a)

OpenStudy (solomonzelman):

(this is also true when your "a" is 5)

OpenStudy (anonymous):

so its √25 and √15 ?

OpenStudy (anonymous):

lskjncxzx square root of 25 and square root of 15?

OpenStudy (solomonzelman):

yes, sqrt(25)-sqrt(15)

OpenStudy (solomonzelman):

but sqrt(25) = 5

OpenStudy (solomonzelman):

so 5-sqrt(13)

OpenStudy (anonymous):

where do u get square root 13 from?

OpenStudy (solomonzelman):

not 13, 15. it is a typo

OpenStudy (solomonzelman):

5-sqrt(15)... like this

OpenStudy (anonymous):

thats the answer?

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

oh o: thank you !!